OpenStudy (anonymous):

I Will Medal And Fan! I need explanation AND answer! Evaluate the following. 12 divided 4 + 9 - 3 x 4

OpenStudy (phi):

have you learned order of operations?

OpenStudy (anonymous):

yes

OpenStudy (idku):

\(12\div 4+9-3\times 4\) like this? (just wanna verify)

OpenStudy (anonymous):

yes

OpenStudy (phi):

you should remember divide first before add or subtract start on the left side and do 12 divided by 4 first. what do you get ?

OpenStudy (anonymous):

3

OpenStudy (phi):

yes, so replace \(12 \div 4\) with 3 what do you have now?

OpenStudy (anonymous):

3 + 9 - 3 x 4

OpenStudy (phi):

ok, now look for any multiply (or divide) if you find one, do it.

OpenStudy (anonymous):

3 x 4 = 12

OpenStudy (phi):

ok, so replace 3*4 with 12 what do we have now?

OpenStudy (anonymous):

3 + 9 - 3 and idk what to do with the 12

OpenStudy (phi):

you had 3+9 - 3*4 you do the multiply

OpenStudy (anonymous):

i did. its 12

OpenStudy (phi):

yes, and how do write the new expression (it has to have a 12 in it)

OpenStudy (anonymous):

idk then

OpenStudy (phi):

3+9 - 3*4 replace 3*4 with 12 (they mean the same thing)

OpenStudy (anonymous):

3 + 9 - 12

OpenStudy (phi):

ok, now with just add subtract, work left to right first do 3+9

OpenStudy (anonymous):

the answer would be 0

OpenStudy (phi):

yes

OpenStudy (anonymous):

3 + 9 = 12, 12 - 12 = 0

OpenStudy (anonymous):

i've got a feeling this is somehow wrong but oh well

OpenStudy (anonymous):

its never been 0

OpenStudy (phi):

if this is the question \[ 12\div 4+9-3\times 4 \] then 0 is the answer

OpenStudy (phi):

in math, if you follow the rules carefully, you get the answer.

OpenStudy (anonymous):

Let's simplify step-by-step. 12 4 +9−3x(4) =3+9+−12x Combine Like Terms: =3+9+−12x =(−12x)+(3+9) =−12x+12 Answer: =−12x+12